Project

General

Profile

20160315MEMO » History » Version 2

Boyd, Steven, 15 March 2016 14:33

1 1 Boyd, Steven
h1. Operations report to MEMO
2
3
* Data-taking
4
** LiH absorber was installed last week with no problems and we are now taking data smoothly
5
** pressure in decay solenoid fridge is still oscillating. Mike/Mark are aware of it.
6
** we've negotiated an increase in ISIS beamloss trip levels to 2.5 V.ms. This should allow
7
us to run at 2.0 rather than 1.3 (which is what we've been averaging) and help with the data rate.
8 2 Boyd, Steven
** Will continue taking LiH data to the end of the user cycle - there is a request from Melissa to take
9
data with He is SSU and air in SSD. However, LiH has priority.
10 1 Boyd, Steven
11
* User cycle 2016/01
12
** Next user cycle starts on April 12th
13
** We've decided that this will mostly be taken up by Hall activities (QP/QD work, realignment work, etc)
14
** We do want to bolster the zero absorber data-set though, and will use two (possibly three) weekends in
15
April to do so.
16
** I'm putting staffing in place for this - and will send out a shift email soon. 
17
18
* Cycle 2016/02 : (June/July)
19
** What we do here very much depends on what is ready
20
** The current plan is to take energy loss data with LiH (or vacuum) and magnetic fields (SSU (E-C-E) + FC + SSD (E-C-E))
21
** We need to understand if (and when) we power M2 in SSD for a cooling measurement